# Quantum plasmonics model of refractive index sensing using photon correlations

### v1

@@ I Introduction
Amongst the quantum effects in a hybrid plasmonic system that have been investigated for plasmonic sensing is the coherent exciton-plasmon dynamics [13, 14] and Rabi splitting [15, 16], where theoretical models were used to show some promising results. However, an interesting quantum effect not yet explored for quantum plasmonic sensing is Fano interference [17, 18]. This is an interference phenomenon that occurs when two or more resonators interact. It is not exclusive to quantum systems [19, 20, 17, 18] and is also actively being explored for classical plasmonic sensing [21, 22, 23]. In the quantum regime, however, several authors have shown that Fano interference is exhibited by hybrid systems of single, or multiple QDs, and single plasmonic nanostructures, such as MNPs and nanocavities of different shapes and sizes [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34].
@@
An intriguing property of these hybrid systems is their ability to support both Poissonian and sub-Poissonian photon statistics in the scattered light, induced by the Fano interference between the resonance of the MNP, for instance, and that of the QD [25, 27, 28, 29]. The interaction between the broad linewidth of the localized surface plasmon resonance (LSPR) of the MNP supporting a continuum of energy and the narrow absorption, or emission, linewidth of the QD supporting a discrete energy gives rise to a Fano resonance and a modification of the quantum statistics of the scattered light [19, 25]. The intensity–intensity correlation function, $g^{(2)}(\tau)$, is often used to characterize the photon statistics of the light scattered by the MNP component of the hybrid system [25], or by the entire hybrid system [27]. A rise in $g^{(2)}(\tau)$ with time ($\tau>0$) indicates photon antibunching, i.e., the tendency of the scattered photons to arrive at a photodetector in a near-regular pattern, while the converse indicates photon bunching [35, 36, 37, 38]. In Refs. [28, 25], it was shown that constructive Fano interference, which occurs within the Fano peak, is responsible for interference-induced photon antibunching and sub-Poissonian statistics, while interference-induced photon bunching and super-Poissonian statistics occurs within the Fano dip due to destructive Fano interference. A question then naturally arises about whether the antibunching with sub-Poissonian statistics found within the Fano peak could possibly be exploited to reduce noise and gain an advantage in sensing.
As a fundamental property of single-photon sources [37, 39, 40, 41], the antibunching effect could potentially be utilized to minimize the shot noise component of the overall noise [42, 43] and improve the resolution of a plasmonic sensor operating at driving wavelengths within the Fano spectrum. Previous quantum studies have investigated the use of plasmonic systems [44] or hybrid emitter-MNP systems displaying Fano effects in designing metamaterials [34] and plasmonic lasers [45, 46]. In quantum information science, plasmonic systems have been used to improve the efficiency of single-photon detectors [10, 11], while emitters such as QDs have proven to be reliable single-photon sources [40, 41]. On the other hand, some authors have demonstrated the generation of non-classical light in a photonic device via Fano interference [20], and remarkably the use of $g^{(2)}(0)$ [47, 12] as well as other higher-order intensity correlations [48, 49] to improve image resolution in confocal microscopy and quantum imaging. Thus, we hypothesize that in a Fano-based quantum plasmonic sensor, the intensity fluctuation in the scattered light could be minimized below the shot noise via photon antibunching, since it depends on $g^{(2)}(0)$ as well as on the scattered intensity, according to photodetection theory [35], thus providing an enhancement in the sensor’s resolution. At the same time, the use of $g^{(2)}(0)$ as a novel sensor signal itself may provide a different route to achieving an enhancement in the sensor’s resolution.
@@
In this work, we investigate the above-mentioned enhancement possibilities by proposing a theoretical model of an emitter-MNP quantum plasmonic sensor that makes use of the scattered intensity of the constructive Fano interference and the second-order correlation function at zero-delay time, $g^{(2)}(0)$, within the Fano spectrum, for refractive index sensing. By calculating the sensitivity and the resolution of the sensor based on the scattered intensity and $g^{(2)}(0)$, a comparison is made between their sensing performance at driving wavelengths within the plasmon and Fano spectra. Although the intensity has been thoroughly investigated for refractive index sensing [50, 51, 52, 53, 1, 22] and applied to biosensing [54, 55], the comparisons between it and $g^{(2)}(0)$ drawn here are necessary to investigate the possibility of an enhancement in each of the sensing parameters—refractive index sensitivity and resolution—predicted by our model.
In addition, we show that when $g^{(2)}(0)$ is utilized as a sensing signal, the signal fluctuations depend on both $g^{(2)}(0)$ and other higher-order correlations up to the fourth order, as well as on the scattered intensity. Furthermore, we find that the mean photocount is inherently low in the emitter-MNP quantum system, inhibiting the role of interference-induced photon antibunching in minimizing the sensor’s shot noise when the intensity is used as a signal. However, we identify a regime in which a quantum enhancement can be achieved in the sensing resolution, requiring an increase in the mean photocount without losing the antibunching effect.
@@ II Model geometry and Input parameters
The system we consider is shown in Fig. 1(a). It consists of an emitter—which we model as a core-shell QD as an example—held by a linker attached to a spherical MNP. The system is embedded in a medium and supported by a substrate. In the configuration shown, referred to as the longitudinal coupling (LC) configuration [27, 25, 24], the electric dipole moments of the MNP and QD are both parallel to the MNP-QD axis in the direction of polarization of the incident field, i.e., $\bm{\chi}=\chi\mathbf{\hat{e}}_{z}$ and $\bm{\mu}=\mu\mathbf{\hat{e}}_{z}$, respectively. We will consider only the LC configuration, since it has been reported as the optimal configuration for atomic cooperativity between the MNP and the QD, as well as for interference-induced effects in the hybrid system [25, 27, 28, 24], though our model can also be applied to the transverse coupling case. The parameter $S_{\beta}$ is the coupling factor of the MNP dipole moment, $\bm{\chi}$, to the image dipole formed in the substrate (perpendicular coupling: $\beta=z,S_{\beta}=2$, parallel coupling: $\beta=x,y,S_{\beta}=1$[56], and $S_{\alpha}$ is the coupling factor of the QD to the MNP (longitudinal coupling: $\alpha=z,S_{\alpha}=2$, transverse coupling: $\alpha=x,y,S_{\alpha}=-1$[25, 24]. Here, we take $S_{\alpha}=S_{\beta}=2$.
@@
The QD is chosen to be a CdTe/CdS core-shell semiconductor nanocrystal whose emission spectrum is tunable within the visible region, as discussed in Ref. [57]. The core-shell radius of the QD is denoted as $a$, where $a=r_{c}+t_{s}$, with $r_{c}$ being the core radius, and $t_{s}$ the shell thickness. The QD is modeled as a two-level system with ground and excited states denoted as |0|0$\rangle$| 0 ⟩ and |1|1$\rangle$| 1 ⟩, respectively. Here, the dipole dephasing rate of the QD is ignored and the non-radiative decay rate of the QD is assumed to be negligible. The latter is justified by their high intrinsic quantum yields [57, 40], while the former is valid within the semi-classical, weak-field limit [25]. The refractive index, free-space radiative decay rate, and emission frequency of the QD are denoted as $n_{d}$, $\gamma_{ex}$, and $\omega_{ex}$, respectively. The values of $r_{c},t_{s},\omega_{ex}$, and $\gamma_{ex}$ were taken from Ref. [57], while the value of $n_{d}$ is taken from Ref. [13]. In Fig. 1(a), the surface-to-surface distance between the MNP and the QD, which we denote as $l$, is occupied by a linker, such that $l\geq 3$ nm, within which the dipole–dipole coupling is valid [27], the atomic cooperativity is high [25], and charge transfer between the MNP and the QD does not occur [58].
[…]
@@
In Fig. 1(b), as a basic example, we provide a potentially more accessible experimental setting for the realization of the theoretical model shown in (a) and discussed above, where the driving electric field, $\mathbf{E}(t)$, is the evanescent field produced by total internal reflection. In this case, when the incident angle is greater than the critical angle, the dominant component of the transmitted field in the region above the substrate is in the $\mathbf{\hat{e}}_{z}$-direction. The amplitude of the electric field would then be modified as $E_{0}=t_{p}(2I_{0}/cn_{s}\varepsilon_{0})^{\frac{1}{2}}(n_{s}/n)$, with $t_{p}$ obtained from the Fresnel equations at the boundary, as described in more detail in Ref. [62].
@@ III Equations of motion
Eq. (1) has been written in a frame rotating at the driving frequency, i.e., with respect to the Hamiltonian $\hbar\omega\left(\hat{a}^{\dagger}\hat{a}+\hat{\sigma}^{\dagger}\hat{\sigma}\right)$, where the exciton (ex) raising and lowering operators of the QD, $\hat{\sigma}^{\dagger}=|1\rangle\langle 0|$ and $\hat{\sigma}=|0\rangle\langle 1|$, respectively, and the plasmon (pl) creation and annihilation operators of the MNP, $\hat{a}^{\dagger}$ and $\hat{a}$, respectively, are time-independent operators. The first two terms in Eq. (1) are the free Hamiltonians of the MNP and QD in the rotating frame, respectively. The third term is the Hamiltonian of the dipole–dipole interaction, while the fourth and fifth terms are the driving-field Hamiltonians of the QD and MNP, respectively. In Eq. (1), $\hbar$ is Dirac’s constant, $\Delta_{pl}=\omega_{pl}-\omega$ is the plasmon-driving field detuning frequency, $\Delta_{ex}=\omega_{ex}-\omega$ is the exciton-driving field detuning frequency, $\Omega_{ex}=E_{0}\mu/2\hbar$ is the Rabi frequency, $\mu=er_{0}$ is the magnitude of the dipole moment of the QD, where $e$ is the electronic charge, and $\Omega_{pl}=E_{0}\chi/2\hbar$ is the driving frequency of the MNP. The LSPR of the MNP, denoted as $\omega_{pl}$, the magnitude of the MNP dipole moment, $\chi$, and the coupling rate of the QD to the MNP, denoted here as $g$, in the presence of a substrate, are obtained as (see Supplementary Information, Sections 1, 2, and 4)
@@
In Eq. (3a), $L$ is the geometric factor of a spherical MNP supported by a substrate [56]. The reflectance of the substrate, denoted here as $\mathcal{R}$, is calculated from the quasi-static Fresnel reflection coefficient, $\mathcal{R}=(\varepsilon_{s}-\varepsilon_{b})/(\varepsilon_{s}+\varepsilon_{b})$, as given in Ref. [56], where $\epsilon_{s}=n_{s}^{2}$ is the dielectric constant of the substrate. For a substrate whose refractive index is greater than that of the medium, $f>2$, resulting in a redshift in the LSPR according to Eq. (2a), while both the dipole moment of the MNP and the coupling rate of the QD to the MNP slightly increase, according to Eqs. (2b) and (2c) respectively. In our model, this is the case for a glass substrate. However, a substrate whose refractive index is less than that of the medium will result in a converse effect. In the absence of a substrate, or one closely matched to that of the medium, such as Teflon, we have $f=2$, and Eqs. (2a), (2b), (2c), and (3b) reduce respectively to the previously reported expressions in Refs. [27, 28]. We have therefore extended the theory to the more practical case where the system is supported by a substrate.
@@
are the non-radiative and radiative decay rates of the MNP, respectively, whose derivations are given in Sections 1 and 4 of the Supplementary Information. The non-radiative decay rate of the MNP is obtained from a Lorentzian approximation of its quasi-static dipole polarizability in the presence of a substrate, and the radiative decay rate is treated phenomenologically by calculating it from the Wigner–Weisskopf formula [63].
@@
Details of the derivation are given in Section 4 of the Supplementary Information. The steady-state solution of Eq. (7a) is
@@
Due to the short plasmon lifetime of the MNP dipole mode, i.e., $\gamma_{pl}^{-1}$ (5.75$\sim$ 5.75∼ 5.75 fs) $\ll\gamma_{ex}^{-1}$ (35$\sim$ 35∼ 35 ns), $\langle\overset{.}{\hat{\sigma}}\rangle$ and $\langle\overset{.}{\hat{\sigma}_{z}}\rangle$ have been solved in the adiabatic limit. In the adiabatic approximation [25, 24], Eq. (8) is used to eliminate the dependence of $\langle\overset{.}{\hat{\sigma}}\rangle$ and $\langle\overset{.}{\hat{\sigma}_{z}}\rangle$, respectively, on $\langle\hat{a}\rangle$ to obtain Eqs. (7b) and (7c), as well as the enhanced radiative decay rate of the QD, $\Gamma$, responsible for the Purcell effect, and the modified Rabi frequency, $\Omega$, respectively, as
@@
The detailed derivation of the above analytical solutions is given in Section 4 of the Supplementary Information. They were validated by performing simulations using the open-source quantum simulation toolbox QuTiP [64, 65]. The simulations were performed within a cQED framework where the MNP is treated as a cavity and the QD as a two-level atom [59, 61, 65]. Using number states with up to 101010 photons to truncate the infinite Hilbert space of the MNP field operator, the optical response of the driven atom-cavity system was found to agree with the analytical solution, as shown via the scattered flux observable in Fig. 2(a) (see next section for details). Besides saving computational time, the analytical solutions are more insightful than simulations since they reveal how the QD excited-state population, $\langle\hat{\sigma}^{\dagger}\hat{\sigma}\rangle$, and the coherence, $\langle\hat{\sigma}\rangle$, are modified by the presence of the MNP. As we show in the next section, these modified terms are important in the description of the photon statistics of the light scattered by the system when operating as a nanosensor.

### v2

@@ I Introduction
Amongst the quantum effects in a hybrid plasmonic system that have been investigated for plasmonic sensing is coherent exciton-plasmon dynamics [13, 14] and Rabi splitting [15, 16], where theoretical models show some promising results. However, an interesting quantum effect not yet explored for quantum plasmonic sensing is Fano interference [17, 18] — an interference phenomenon that occurs when two or more resonators interact. Previous studies have shown that Fano interference is exhibited by hybrid systems of single, or multiple QDs, and single plasmonic nanostructures [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Another intriguing property of these hybrid systems is their ability to support both Poissonian and sub-Poissonian photon statistics in the scattered light, induced by the Fano interference between the resonance of the MNP, for instance, and that of the QD [20, 22, 23, 24]. The interaction between the broad linewidth of the localized surface plasmon resonance (LSPR) of the MNP supporting a continuum of energy and the narrow absorption, or emission, linewidth of the QD supporting a discrete energy gives rise to a Fano resonance and a modification of the quantum statistics of the scattered light [30, 20]. The intensity–intensity correlation function, $g^{(2)}(\tau)$, is often used to characterize the photon statistics of the light scattered by the MNP component of the hybrid system [20], or by the entire hybrid system [22]. A rise in $g^{(2)}(\tau)$ with time ($\tau>0$) indicates photon antibunching, i.e., the tendency of the scattered photons to arrive at a photodetector in a near-regular pattern, while the converse indicates photon bunching [31, 32, 33, 34]. In Refs. [23, 20], it was shown that constructive Fano interference, which occurs within the Fano peak, is responsible for interference-induced photon antibunching and sub-Poissonian statistics. A question then naturally arises about whether the antibunching with sub-Poissonian statistics found within the Fano peak could possibly be exploited to reduce noise and gain an advantage in refractive index sensing.
@@
In this work, we seek an answer to the above question by proposing a theoretical model of an emitter-MNP quantum plasmonic sensor that makes use of the scattered intensity of the constructive Fano interference and the second-order correlation function at zero-delay time, $g^{(2)}(0)$, within the Fano spectrum, for refractive index sensing. The use of $g^{(2)}(0)$ and other higher-order intensity correlations has recently been considered for improving image resolution in confocal microscopy and quantum imaging [35, 12, 36, 37], thus providing additional motivation for its study in our context. By calculating the sensitivity and the resolution of the sensor based on the scattered intensity and $g^{(2)}(0)$, a comparison is made between their sensing performance at driving wavelengths within the plasmon and Fano spectra. Although the intensity has been thoroughly investigated for refractive index sensing [38, 39, 40, 41, 1, 42] and applied to biosensing [43, 44], the antibunching effect could potentially minimize the shot noise component of the intensity noise [45, 46] and improve the resolution. Moreover, the comparisons between the intensity and $g^{(2)}(0)$ drawn here are necessary to investigate the possibility of an enhancement in each of their respective sensing parameters—refractive index sensitivity and resolution—predicted by our model.
@@ II Model geometry and Input parameters
The system we consider is shown in Fig. 1(a). It consists of an emitter—which we model as a core-shell QD as an example—held by a linker attached to a spherical MNP. The system is embedded in a background medium and supported by a substrate. In the configuration shown, referred to as the longitudinal coupling (LC) configuration [22, 20, 19], the electric dipole moments of the MNP and QD are both parallel to the MNP-QD axis in the direction of polarization of the incident field, i.e., $\bm{\chi}=\chi\mathbf{\hat{e}}_{z}$ and $\bm{\mu}=\mu\mathbf{\hat{e}}_{z}$, respectively. We will consider only the LC configuration, since it has been reported as the optimal configuration for atomic cooperativity between the MNP and the QD, as well as for interference-induced effects in the hybrid system [20, 22, 23, 19], though our model can also be applied to the transverse coupling case. The parameter $S_{\beta}$ is the coupling factor of the MNP dipole moment, $\bm{\chi}$, to the image dipole formed in the substrate (perpendicular coupling: $\beta=z,S_{\beta}=2$, parallel coupling: $\beta=x,y,S_{\beta}=1$[47], and $S_{\alpha}$ is the coupling factor of the QD to the MNP (longitudinal coupling: $\alpha=z,S_{\alpha}=2$, transverse coupling: $\alpha=x,y,S_{\alpha}=-1$[20, 19]. Here, we take $S_{\alpha}=S_{\beta}=2$.
@@
The QD is chosen to be a CdTe/CdS core-shell semiconductor nanocrystal whose emission spectrum is tunable within the visible region, as discussed in Ref. [48]. The core-shell radius of the QD is denoted as $a$, where $a=r_{c}+t_{s}$, with $r_{c}$ being the core radius, and $t_{s}$ the shell thickness. The QD is modeled as a two-level system with ground and excited states denoted as $|0\rangle$ and $|1\rangle$, respectively. Here, the dipole dephasing rate of the QD is ignored and the non-radiative decay rate of the QD is assumed to be negligible. The latter is justified by their high intrinsic quantum yields [48, 49], while the former is valid within the semi-classical, weak-field limit [20]. The refractive index, free-space radiative decay rate, and emission frequency of the QD are denoted as $n_{d}$, $\gamma_{ex}$, and $\omega_{ex}$, respectively. The values of $r_{c},t_{s},\omega_{ex}$, and $\gamma_{ex}$ were taken from Ref. [48], while the value of $n_{d}$ is taken from Ref. [13]. In Fig. 1(a), the surface-to-surface distance between the MNP and the QD, which we denote as $l$, is occupied by a linker, such that $l\geq 3$ nm, within which the dipole–dipole coupling is valid [22], the atomic cooperativity is high [20], and charge transfer between the MNP and the QD does not occur [50]. As shown in Ref. [23], both Fano interference and antibunching effects are supported by the hybrid system even beyond the dipole-dipole coupling limit.
A gold nanoparticle is used as the model MNP whose radius, denoted as $r$, is large enough such that the local response approximation is valid [20, 22, 29], but small enough to ensure a non-vanishing Fano profile, and subsequently, a sizeable photon antibunching effect, as previously shown in Ref. [22]. The center-to-center distance between the MNP and the QD is denoted as $d$, where $d=r+l+a$. The nanosensor is driven weakly by a $z$-polarized electric field, $\mathbf{E}(t)=E_{0}\cos(\omega t)\mathbf{\hat{e}}_{z}$, with driving frequency $\omega$ and amplitude $E_{0}=(2I_{0}/cn\varepsilon_{0})^{\frac{1}{2}}$, where $I_{0}$ is the incident field intensity, $c$ is the speed of light in vacuum, $n$ is the refractive index of the background medium, and $\varepsilon_{0}$ is the permittivity of free space. The electric field is propagating in the $x$-direction with a wavevector of magnitude $k=\omega n/c$, as shown in Fig. 1(a). We have ignored the spatial dependence of $\mathbf{E}(t)$ since $kx\ll 1$ over the lateral extent of the sensor ($2r$ in the $x$-direction) — the quasi-static (long-wavelength) limit [51, 52, 53, 54]. A weak-driving field ensures that the Fano peak is not suppressed relative to the plasmon resonance peak, since the Fano profile vanishes at strong-driving fields due to the saturation of the QD excited-state population [22, 20].
@@
In Fig. 1(b), as a basic example, we provide a potentially more accessible experimental setting for the realization of the theoretical model shown in (a) and discussed above. It is based on a setup experimentally realized in Ref. [55], thus making our scenario more realistic. Here, the driving electric field, $\mathbf{E}(t)$, is the evanescent field produced by total internal reflection. In this case, when the incident angle is greater than the critical angle, the dominant component of the transmitted field in the region above the substrate is in the $\mathbf{\hat{e}}_{z}$-direction. The amplitude of the electric field would then be modified as $E_{0}=t_{p}(2I_{0}/cn_{s}\varepsilon_{0})^{\frac{1}{2}}(n_{s}/n)$, with $t_{p}$ obtained from the Fresnel equations at the boundary, as described in more detail in Ref. [56].
@@ III Equations of motion
Eq. (1) has been written in a frame rotating at the driving frequency, i.e., with respect to the Hamiltonian $\hbar\omega\left(\hat{a}^{\dagger}\hat{a}+\hat{\sigma}^{\dagger}\hat{\sigma}\right)$, where the exciton (ex) raising and lowering operators of the QD, $\hat{\sigma}^{\dagger}=|1\rangle\langle 0|$ and $\hat{\sigma}=|0\rangle\langle 1|$, respectively, and the plasmon (pl) creation and annihilation operators of the MNP, $\hat{a}^{\dagger}$ and $\hat{a}$, respectively, are time-independent operators. The first two terms in Eq. (1) are the free Hamiltonians of the MNP and QD in the rotating frame, respectively. The third term is the Hamiltonian of the dipole–dipole interaction, while the fourth and fifth terms are the driving-field Hamiltonians of the QD and MNP, respectively. In Eq. (1), $\hbar$ is Dirac’s constant, $\Delta_{pl}=\omega_{pl}-\omega$ is the plasmon-driving field detuning frequency, $\Delta_{ex}=\omega_{ex}-\omega$ is the exciton-driving field detuning frequency, $\Omega_{ex}=E_{0}\mu/2\hbar$ is the Rabi frequency, $\mu=er_{0}$ is the magnitude of the dipole moment of the QD, where $e$ is the electronic charge, and $\Omega_{pl}=E_{0}\chi/2\hbar$ is the driving frequency of the MNP. The LSPR of the MNP, denoted as $\omega_{pl}$, the magnitude of the MNP dipole moment, $\chi$, and the coupling rate of the QD to the MNP, denoted here as $g$, in the presence of a substrate, are obtained as (see Sections 1, 2, and 4 of the Supplemental Material [57])
@@
In Eq. (3a), $L$ is the geometric factor of a spherical MNP supported by a substrate [47]. This factor allows us to generalize previous work to the case where the system is placed on a substrate, which is more physically relevant. The reflectance of the substrate, denoted here as $\mathcal{R}$, is calculated from the quasi-static Fresnel reflection coefficient, $\mathcal{R}=(\varepsilon_{s}-\varepsilon_{b})/(\varepsilon_{s}+\varepsilon_{b})$, as given in Ref. [47], where $\epsilon_{s}=n_{s}^{2}$ is the dielectric constant of the substrate. For a substrate whose refractive index is greater than that of the medium, $f>2$, resulting in a redshift in the LSPR according to Eq. (2a), while both the dipole moment of the MNP and the coupling rate of the QD to the MNP slightly increase, according to Eqs. (2b) and (2c) respectively. In our model, this is the case for a glass substrate. However, a substrate whose refractive index is less than that of the medium will result in a converse effect. In the absence of a substrate, or one closely matched to that of the medium, such as Teflon, we have $f=2$, and Eqs. (2a), (2b), (2c), and (3b) reduce respectively to the previously reported expressions derived in Refs. [22, 23]. We have therefore extended the theory to the more practical case where the system is supported by a substrate, which is relevant to experimental setups.
@@
are the non-radiative and radiative decay rates of the MNP, respectively, whose derivations are given in Sections 1 and 4 of the Supplemental Material [57]. The non-radiative decay rate of the MNP is obtained from a Lorentzian approximation of its quasi-static dipole polarizability in the presence of a substrate, and the radiative decay rate is treated phenomenologically by calculating it from the Wigner–Weisskopf formula [58], which is consistent with a treatment using the radiation reaction field [59].
@@
Details of the derivation are given in Section 4 of the Supplemental Material [57]. The steady-state solution of Eq. (7a) is
@@
Due to the short plasmon lifetime of the MNP dipole mode, i.e., $\gamma_{pl}^{-1}$ ($\sim 5.75$ fs) $\ll\gamma_{ex}^{-1}$ ($\sim 35$ ns), $\langle\overset{.}{\hat{\sigma}}\rangle$ and $\langle\overset{.}{\hat{\sigma}_{z}}\rangle$ have been solved in the adiabatic limit. In the adiabatic approximation [20, 19], Eq. (8) is used to eliminate the dependence of $\langle\overset{.}{\hat{\sigma}}\rangle$ and $\langle\overset{.}{\hat{\sigma}_{z}}\rangle$, respectively, on $\langle\hat{a}\rangle$ to obtain Eqs. (7b) and (7c), as well as the enhanced radiative decay rate of the QD, $\Gamma$, responsible for the Purcell effect, and the modified Rabi frequency, $\Omega$, respectively, as
@@
The detailed derivation of the above analytical solutions is given in Section 4 of the Supplemental Material [57]. They were validated by performing simulations using the open-source quantum simulation toolbox QuTiP [60, 61]. The simulations were performed within a cQED framework where the MNP is treated as a cavity and the QD as a two-level atom [51, 53, 61]. Using number states with up to $10$ photons to truncate the infinite Hilbert space of the MNP field operator, the optical response of the driven atom-cavity system was found to agree with the analytical solution, as shown via the scattered flux observable in Fig. 2(a) (see next section for details). Besides saving computational time, the analytical solutions are more insightful than simulations since they reveal how the QD excited-state population, $\langle\hat{\sigma}^{\dagger}\hat{\sigma}\rangle$, and the coherence, $\langle\hat{\sigma}\rangle$, are modified by the presence of the MNP. As we show in the next section, these modified terms are important in the description of the photon statistics of the light scattered by the system when operating as a nanosensor.